# The Automorphism Equivariant Hitchin Index

### v1

@@ 3 Localization
We collect some generalities on localization before proving Theorem 1.2. Let $H$ be a finite cyclic group with generator $h.$ Let $(X,\omega)$ be a Kähler manifold with a holomorphic $\mathbb{T}\times H$-action, for which the $\mathbb{T}$-action is meromorphic in the sense of [68], i.e. assume Assumption $2.15$ in [68, §2.2] is satisfied. Let $L\rightarrow X$ be a holomorphic $\mathbb{T}\times H$-equivariant line bundle of curvature $-i\omega$. Then $L\rightarrow X$ is a quantum line bundle for $(X,\omega)$ in the sense of geometric quantization [47, 58]. For all $q\in\mathbb{Z}_{+}$ and $m\in\mathbb{Z}$ denote by $H^{q}_{m}(X,L)$ the $t^{m}$-weight sub-space of $H^{q}(X,L)$. Define
$\chi_{\mathbb{T}}(X,L,h)(t)=\sum_{m\in\mathbb{Z}}\sum_{q\geq 0}(-1)^{q}\operatorname{tr}(h:H_{m}^{q}(X,L)\rightarrow H_{m}^{q}(X,L))\cdot t^{m}.$ (72)
The meromorphicity condition on the $\mathbb{T}$-action on $X$ implies that the connected components of $X^{\mathbb{T}}$ are compact smooth complex manifolds. Write $A:=\pi_{0}(X^{\mathbb{T}\times H})$ and denote by $\{X_{a}\}_{a\in A}$ the components of $X^{\mathbb{T}\times H}$. Fix $a\in A$. Let $T_{a}X$ denote the restriction of $TX$ to $X_{a}$. For every $\mathbb{T}\times H$-equivariant vector bundle $W\rightarrow X_{a}$ denote by $W^{\pm}$ the $\pm$-$\mathbb{T}$-weight subbundle of $W$, and define $\nu_{a}=\operatorname{rank}(T_{a}X^{-}).$ Recall the notation $s$ for the equivariant sum of all symmetric powers. Define the equivariant class
$\omega(W)=s((W^{+})^{*})\cdot s(W^{-})\cdot\det(W^{-}).$ (73)
Denote by $N^{*}(X_{a})$ the conormal of $X_{a}$ as a submanifold of $X^{\mathbb{T}}$. Denote by $\lambda=\sum_{l\geq 0}(-1)^{l}\lambda^{l}$ the equivariant operation of taking the alternating sum of exterior powers.
@@
$\begin{split}&\chi_{\mathbb{T}}(X,L,h)=\sum_{a\in A}(-1)^{\nu_{a}}\int_{X_{a}}\operatorname{ch}\left(\frac{\omega(T_{a}X)\cdot L_{\mid X_{a}}}{\lambda\left(N^{*}(X_{a})\right)}\right)\operatorname{Td}X_{a}.\end{split}$ (74)
Proposition 3.1 follows essentially from a combination of a $\mathbb{T}$-localization result of Wu [68] with the Atiyah-Bott fixed point formula [14], also known as the Lefshetz fixed-point formula. In [68], Wu considers the case $H=\langle 1\rangle$. In this case $\{X_{a}\}_{a\in A}$ is equal to the set of components of $X^{\mathbb{T}}$ and in [68, §4.2] the following is proven.
@@
$\begin{split}&\sum_{\begin{subarray}{c}m\in\mathbb{Z},\\ q\geq 0\end{subarray}}(-1)^{q}H^{q}_{m}(X,L)\cdot t^{m}=\sum_{\begin{subarray}{c}m\in\mathbb{Z},\\ q\geq 0\end{subarray}}\sum_{a\in A}(-1)^{\nu_{a}+q}H_{m}^{q}(X_{a},\omega(T_{a}X)\cdot L_{\mid X_{a}})\cdot t^{m}.\end{split}$ (75)
We now recall the Atiyah-Bott Lefshetz formula [14]. Let $Z$ be a compact complex manifold with a holomorphic $H$-action. Let $W\rightarrow Z$ be an $H$-equivariant holomorphic vector bundle. Consider the equivariant index
@@
Let $Z^{H}$ denote the fixed locus. Write $B:=\pi_{0}(Z^{H})$ and denote by $\{Z_{b}\}_{b\in B}$ the components of $Z^{H}$. For each component $Z_{b}$, denote by $N^{*}(Z_{b})$ the conormal of $Z_{b}\subset Z$. The components are of course compact and smooth.
@@
The proof is a simple computation which combines (75) with (76). The key point is that the identity (75) can be generalized to an identity in the $K$-theory ring of $\mathbb{T}\times H$-representations. ∎
We now recall Nielsens’ localization result [52]. Assume now that $H=\mu_{p}$ and that $Z$ is a smooth projective $\mu_{p}$-variety. Let $K_{\mu_{p}}^{0}(Z)$ denote the $K$-theory ring of $\mu_{p}$-equivariant holomorphic vector bundles. Let $Z^{\mu_{p}}$ denote the fixed variety. Recall that $\mathbb{Z}[\mu_{p}]$ denotes the representation ring of $\mu_{p}$. Recall the multiplicatively closed subset $S\subset\mathbb{Z}[\mu_{p}]$ defined in Section 1.2. We can regard $S$ as a multiplicatively closed subset of $K^{0}_{\mu_{p}}(Z)$ and of $K^{0}_{\mu_{p}}(Z^{\mu_{p}})$.
@@
We now analyze $K_{\mu_{p}}^{0}(Z^{\mu_{p}})$ in more detail. We have canonical isomorphisms
@@
For $j\in\{0,...,p-1\}$ write $\pi_{j}$ for the projection onto the $j$’th summand in (78). We introduce the special notation $W_{0}=\pi_{0}(W).$ We denote the universal localization homomorphism by $L:K^{0}_{\mu_{p}}(Z^{\mu_{p}})\rightarrow S^{-1}\cdot K^{0}_{\mu_{p}}(Z^{\mu_{p}}).$ Recall that the kernel of $L$ is generated by all elements $r$ such that $s\cdot r=0$ for some $s\in S$. Let $Y$ be a formal variable. Recall that $\mathbb{Z}[\mu_{p}]\cong\mathbb{Z}[Y]/(Y^{P}-1)$ and that $Y^{P}-1=(Y-1)\Phi_{p}(Y)$ is the factorization of $Y^{p}-1$ into irreducible factors. As $(1-\hat{\zeta})\in S$, it is therefore clear that we have $\ker(L)=K^{0}(Z^{\mu_{p}})\otimes_{\mathbb{Z}}(\Phi_{p}(\hat{\zeta})).$ Moreover, the following sequence of abelian groups is split exact
@@
and a complement to $\ker(L)$ is given by the module generated by all classes of the form $W-W_{0}\cdot\Phi_{p}(\hat{\zeta})$, where $W$ ranges over all equivariant bundles. The following lemma is an elementary consequence of these facts.
@@
$V-V_{0}\cdot\Phi_{p}(\hat{\zeta})=W-W_{0}\cdot\Phi_{p}(\hat{\zeta}).$ (80)
Recall now the setup and the notation introduced in Section 1.2. For each $l\in\{0,...,p-1\}$, let $\pi_{l}$ be the projection onto the $l$’th summand of $R_{i,j}$ in (5).
@@
It will be seen in the proof of Theorem 1.2 that the $\tilde{a}_{i,j,l}$ and the $\tilde{b}_{i,j,l}$ are in fact positive integers. Before given the proof of Theorem 1.2 we state the following Lemma.
@@
Write $a=\sum_{i=0}^{\infty}a_{i}t^{-1}$ and $b=\sum_{i=0}^{\infty}b_{i}t^{-i}$ with $a_{i},b_{i}\in R$ for all non-negative integers $i$. Without loss of generality, we may assume $a_{0}\not=0$ and $b_{0}\not=0$. We have
@@
In particular, we have that $a_{0}^{p}=b_{0}^{p}$. For each positive integer $l$, let $R^{l}$ denote the degree $2l$ part of $R$ with respect to the cohomological grading. Write $a_{0}=\sum_{l=0}^{d}a_{0,l}$ and $b_{0}=\sum_{l=0}^{d}b_{0,l}$ with $a_{0,l},b_{0,l}\in R^{l}$ for $l=0,...,d$. Then we have $a_{0,0}=b_{0,0}$ by assumption, and we see that
@@
For each $l=1,..,d$ we get by taking the degree $l$ part of (85) that
@@
Now notice that any element of $R$ whose projection onto $R^{0}$ is non-zero is not a zero divisor in $R$. In particular $a_{0,0}^{p}=b_{0,0}^{p}$ is not a zero divisor, and therefore we get by induction from (86) that $a_{0,l}=b_{0,l}$ for $l=0,...,d$. Thus $a_{0}=b_{0}$. From (84) we deduce that for all non-negative integers $i$ we have that
@@
By induction, and the fact that $a_{0}^{p}$ and $b_{0}^{p}$ are not zero divisors, we obtain from (87) that $a_{i}=b_{i}$ for all non-negative integers $i$. Thus $a=b$. ∎
Proof of Theorem 1.2.
Fix $i\in\mathcal{I}$. Write $\alpha_{i}$ (resp. $\tilde{\alpha}_{i}$) for the equivariant first Chern class of the restriction of $\mathcal{L}$ (resp. ~$\tilde{\mathcal{L}}_{i}$) to $\mathcal{M}_{i}$ (resp. ~$\tilde{\mathcal{M}}_{i}$). Then we have that
$p\cdot\tilde{\alpha}_{i}=(\varXi_{i}^{-1})^{*}(\alpha_{i}).$ (88)
For the restriction $\varXi_{i}:\mathcal{N}\cap\mathcal{M}_{i}\rightarrow\tilde{\mathcal{N}}_{i}$, the identity (88) was in fact proven in the article [3], and the proof therein can be generalized to the isomorphism of $\mathbb{T}$-varieties $\varXi_{i}:\mathcal{M}_{i}\rightarrow\tilde{\mathcal{M}}_{i}$. We now turn to the proof of equation (9). The bundle $\mathcal{L}\rightarrow\mathcal{M}$ satisfies the conditions of Proposition 3.1 with $H=\langle f\rangle$ and the $\mathbb{T}\times\langle f\rangle$-equivariant structure constructed in Section 2. Thus we can apply Proposition 3.1. Let $i\in\mathcal{I}$ and $j\in\mathcal{J}_{i}$. To ease notation, we will write $\mathcal{M}_{i,j}^{\mathbb{T}}=\mathcal{M}_{i,j}$ and we will write $T_{i,j}\mathcal{M}$ (resp. $T_{i,j}\mathcal{M}_{i}$) for the restriction of $T\mathcal{M}$ (resp. $T\mathcal{M}_{i}$) to $\mathcal{M}_{i,j}$. Let $N^{*}_{i,j}$ be the conormal of $\mathcal{M}_{i,j}\subset\mathcal{M}^{\mathbb{T}}$. By Proposition 3.1, it will be sufficient to prove that for all $i\in\mathcal{I}$ and $j\in\mathcal{J}_{i}$ we have that
$(-1)^{\nu_{i,j}}\omega(T_{i,j}\mathcal{M})\cdot\lambda\left(N^{*}_{i,j}\right)^{-1}=(-1)^{\tilde{\nu}_{i,j}}\varXi_{i}^{*}(\tilde{\chi}_{i,j}).$ (89)
Write $\mathbb{U}=\operatorname{End}_{0}(\operatorname{\mathbbm{E}})$, and define the equivariant class
@@
This can be seen for instance by considering a standard long exact sequence for hypercohomology groups of two-term chain complexes (see [36, §2.2.3]). Consider the natural projection onto the first factor $\pi^{\prime}:\mathcal{M}_{i,j}\times X^{f}\rightarrow\mathcal{M}_{i,j}$. Note that we have $(\mathcal{M}_{i,j}\times X)^{f}=\mathcal{M}_{i,j}\times X^{f}$. Consider the inclusion $\iota:\mathcal{M}_{i,j}\times X^{f}\rightarrow\mathcal{M}_{i,j}\times X$. Clearly $\pi\circ\iota=\pi^{\prime}.$ By applying Nielsen’s localization theorem to the right hand side of quation (91), we obtain the following identity in $S^{-1}\cdot K^{0}_{\mathbb{T}\times\mu_{p}}(\mathcal{M}_{i,j})$
@@
However, the $\mathbb{T}\times\mu_{p}$-equivariant $K$-theory operation $s$ of taking the sum of all symmetric powers is not well-defined in $S^{-1}\cdot K^{0}_{\mathbb{T}\times\mu_{p}}(\mathcal{M}_{i,j})$, and we need to extract an identity in the ring $K_{\mathbb{T}\times\mu_{p}}^{0}(\mathcal{M}_{i,j})$. For all $u=1,...,p-1$, we have the following identity in $S^{-1}\cdot\mathbb{Z}[\mu_{p}]$

### v2

@@ 2 Localization
We collect some generalities on localization before proving Theorem LABEL:ThmMain. Let $(M,\omega)$ be a Kähler manifold with a holomorphic $\mathbb{T}\times\mu_{p}$-action, for which the $\mathbb{T}$-action is meromorphic in the sense of [79], i.e. assume Assumption $2.15$ in [79, §2.2] is satisfied. This is satisfied if $M$ is a semi-projective variety in the sense of [23]. Let $h:M\rightarrow M$ be a generator of the $\mu_{p}$-action. Let $L\rightarrow M$ be a holomorphic $\mathbb{T}\times\mu_{p}$-equivariant line bundle of curvature $i\omega$. Then $L\rightarrow M$ is a quantum line bundle for $(M,\omega)$ in the sense of geometric quantization [56, 69]. For all $q\in\mathbb{Z}_{+}$ and $m\in\mathbb{Z}$ denote by $H^{q}_{m}(M,L)$ the $t^{m}$-weight sub-space of $H^{q}(M,L)$. Finite dimensionality of these weight sub-spaces follows from the results of [79]. Define
$\chi_{\mathbb{T}}(M,L,h)(t)=\sum_{m\in\mathbb{Z}}\sum_{q\geq 0}(-1)^{q}\operatorname{tr}(h:H_{m}^{q}(M,L)\rightarrow H_{m}^{q}(M,L))\cdot t^{m}.$ (75)
Proposition 2.1 below gives a general localization formula for $\chi_{\mathbb{T}}(M,L,h)$ which we will apply to the automorphism equivariant Hitchin index as defined in (LABEL:eq:index). To state Proposition 2.1, we first introduce some notation.
The meromorphicity condition on the $\mathbb{T}$-action on $M$ implies that the connected components of $M^{\mathbb{T}}$ are compact smooth complex manifolds. For our purpose, we can and will assume each of these components is in fact a smooth projective variety. Write $A:=\pi_{0}(M^{\mathbb{T}\times\mu_{p}})$ and denote by $\{M_{a}\}_{a\in A}$ the components of $M^{\mathbb{T}\times\mu_{p}}$. Fix $a\in A$. Let $T_{a}M$ denote the restriction of $TM$ to $M_{a}$. For every $\mathbb{T}\times\mu_{p}$-equivariant vector bundle $W\rightarrow M_{a}$ denote by $W^{\pm}$ the $\pm$-$\mathbb{T}$-weight subbundle of $W$, and define $\nu_{a}=\operatorname{rank}(T_{a}M^{-}).$ Recall the notation $s$ for the equivariant sum of all symmetric powers. Recall that for any algebraic group $H$, we denote by $K_{H}^{0}$ the functor that assigns the K-theory ring of $H$-equivariant algebraic vector bundles to a smooth projective $H$-variety. Define the equivariant class
$\omega(W)=s((W^{+})^{*})\cdot s(W^{-})\cdot\det(W^{-})\in K^{0}_{\mathbb{T}\times\mu_{p}}(M_{a})[[t^{-1}]].$ (76)
The equivariant operation $\Omega$ introduced in (LABEL:def:Omega) in Section LABEL:sec:introlocalization play a key role in Proposition 2.1 and we now justify why it is well-defined. Denote by $\lambda=\sum_{l\geq 0}(-1)^{l}\lambda^{l}$ the equivariant operation of taking the alternating sum of exterior powers. Let $W^{\mathbb{T}}\subset W$ (resp. $W^{\mathbb{T}\times\mu_{p}}\subset W)$ denote the sub-bundle fixed by $\mathbb{T}$ (resp. $\mathbb{T}\times\mu_{p}$). Recall from Section LABEL:sec:introlocalization the multplicatively closed subset $S\subset\mathbb{Z}[\mu_{p}]$ generated by the elements $1-\hat{\zeta^{l}}$ where $l$ runs through the set $\{1,...,p-1\}.$ For any smooth projective $H$-variety $Z^{\prime}$ we observe that $K_{\mu_{p}}(Z^{\prime})$ is an algebra over the ring $\mathbb{Z}[\mu_{p}]$, and we can therefore regard $S$ as a multiplicatively closed subset of $K^{0}_{\mu_{p}}(Z^{\prime})$. To justify why $\Omega(W)$ is well-defined, we must justify the invertibility of the image of the class $\lambda((W^{\mathbb{T}}/W^{\mathbb{T}\times\mu_{p}})^{*})$ under the natural homomorphism of rings $K^{0}_{\mu_{p}}(M_{a})\rightarrow S^{-1}\cdot K^{0}_{\mu_{p}}(M_{a})\otimes\mathbbm{Q}$. This follows from Proposition 2.2 below. We now make the simple observation that we may rewrite $\Omega(W)$ in terms of $\omega(W)$ as follows
$\Omega(W)=\frac{\omega(W)}{\lambda((W^{\mathbb{T}}/W^{\mathbb{T}\times\mu_{p}})^{*})}\in S^{-1}\cdot K^{0}_{\mathbb{T}\times\mu_{p}}(M_{a})\otimes_{\mathbb{Z}}\mathbbm{Q}[[t^{-1}]].$ (77)
Further, we observe that $\Omega$ is a homomorphism that maps sums to products.
@@
$\begin{split}&\chi_{\mathbb{T}}(M,L,h)=\sum_{a\in A}(-1)^{\nu_{a}}\int_{M_{a}}\operatorname{ch}\left(\Omega(T_{a}M)\cdot L_{\mid M_{a}}\right)\operatorname{Td}M_{a}.\end{split}$ (78)
Proposition 2.1 follows essentially from a combination of a $\mathbb{T}$-localization result of Wu [79] with the Atiyah-Bott fixed point formula [19], also known as the Lefshetz fixed-point formula. In [79], Wu considers the case where $\mu_{p}$ act trivially on $M$. In this case $\{M_{a}\}_{a\in A}$ is equal to the set of components of $M^{\mathbb{T}}$ and in [79, §4.2] the following is proven.
@@
$\begin{split}&\sum_{\begin{subarray}{c}m\in\mathbb{Z},\\ q\geq 0\end{subarray}}(-1)^{q}H^{q}_{m}(M,L)\cdot t^{m}=\sum_{\begin{subarray}{c}m\in\mathbb{Z},\\ q\geq 0\end{subarray}}\sum_{a\in A}(-1)^{\nu_{a}+q}H_{m}^{q}(M_{a},\omega(T_{a};)\cdot L_{\mid M_{a}})\cdot t^{m}.\end{split}$ (79)
We now recall the Atiyah-Bott Lefshetz formula [19]. Let $Z$ be a smooth projective complex variety with an algebraic $\mu_{p}$-action generated by an isomorphism $h:Z\rightarrow Z$. Let $W\rightarrow Z$ be a $\mu_{p}$-equivariant algebraic vector bundle. Consider the equivariant index
@@
Let $Z^{\mu_{p}}$ denote the fixed locus. Write $B:=\pi_{0}(Z^{\mu_{p}})$ and denote by $\{Z_{b}\}_{b\in B}$ the components of $Z^{\mu_{p}}$. For each component $Z_{b}$, denote by $N^{*}(Z_{b})$ the conormal of $Z_{b}\subset Z$. The components are of course smooth and projective. Fix $b\in B$. We have natural isomorphisms
$K^{0}_{\mu_{p}}(Z_{b})\cong\mathbb{Z}[\mu_{p}]\otimes_{\mathbb{Z}}K^{0}(Z_{b})\cong\bigoplus_{l=0}^{p-1}\hat{\zeta}^{l}\cdot K^{0}(Z_{b}),$ (80)
the leftmost of which is an isomorphism of $\mathbb{Z}[\mu_{p}]$-algebras and the rightmost of which is an isomorphism of abelian groups. For each $l\in\{0,...,p-1\}$, we denote by $\pi_{l}:K^{0}_{\mu_{p}}(Z_{b})\rightarrow K^{0}(Z_{b})$ the projection onto the coefficient of $\hat{\zeta}^{l}$. Similarly, for an equivariant algebraic vector bundle $V\rightarrow Z_{b}$, we also use the notation $\pi_{l}(V)$ to denote the algebraic vector sub-bundle of $V\rightarrow Z_{b}$ on which $\mu_{p}$ act with weight $\hat{\zeta}^{l}$. Recall the multiplicatively closed subset $S\subset\mathbb{Z}[\mu_{p}]$ generated by the elements $1-\hat{\zeta}^{l}$, where $l$ runs through the set $\{1,...,p-1\}.$
Proposition 2.2.
Let $V\rightarrow Z_{b}$ be a $\mu_{p}$-equivariant algebraic vector bundle. Then $\lambda(V/V^{\mu_{p}}))$ is invertible in $S^{-1}\cdot K^{0}_{\mu_{p}}(Z)\otimes\mathbbm{Q}$. Moreover, the equivariant rank $\operatorname{rank}(\lambda(V/V^{\mu_{p}}))$ is an element of $S$, and if we set $r_{l}=\operatorname{rank}(\pi_{l}(V))$, then the following two formuale holds
$\begin{split}&\operatorname{rank}(\lambda(V/V^{\mu_{p}}))=\prod_{l=1}^{p-1}(1-\hat{\zeta}^{l})^{r_{l}},\\ &\lambda(V/V^{\mu_{p}})^{-1}=\operatorname{rank}(\lambda(V/V^{\mu_{p}}))^{-1}\sum_{a=0}^{\dim(Z_{b})}\left(\frac{(\operatorname{rank}(\lambda(V/V^{\mu_{p}}))-\lambda(V/V^{\mu_{p}})}{\operatorname{rank}(\lambda(V/V^{\mu_{p}}))}\right)^{a}.\end{split}$ (81)
The first part of this proposition is often stated in the case where $V=T^{*}_{\mid Z_{b}}Z$, so that $V/V^{\mu_{p}}$ is the conormal of $Z_{b}\hookrightarrow Z$. For the sake of completeness, we present a proof.
Proof.
Let $A(Z_{b})$ be the Chow ring of $Z_{b}$. By [34, Example 15.2.16]) the Chern character induces an isomorphism $\operatorname{ch}:K^{0}(Z_{b})\otimes_{\mathbb{Z}}\mathbbm{Q}\overset{\sim}{\rightarrow}A(Z_{b})\otimes_{\mathbb{Z}}\mathbbm{Q}.$ Clearly, in our setting, this implies that the equivariant Chern character induces an isomorphism
$\operatorname{ch}:S^{-1}\cdot K_{\mu_{p}}^{0}(Z_{b})\otimes_{\mathbb{Z}}\mathbbm{Q}\overset{\sim}{\rightarrow}S^{-1}\cdot\mathbb{Z}[\mu_{p}]\otimes_{\mathbb{Z}}A(Z_{b})\otimes_{\mathbb{Z}}\mathbbm{Q}.$ (82)
We will refer to the target ring of this isomorphism as the equivariant Chow ring. Define algebraic vector bundles $V_{l}$ where $l$ runs through the set $\{1,...,p-1\}$ by $\pi_{l}(V)=V_{l}$. As $\lambda$ is a homomorphism we have that
$\lambda(V/V^{\mu_{p}})=\prod_{l=1}^{p-1}\lambda(\hat{\zeta}^{l}\cdot V_{l}).$ (83)
It is enough to argue that for each $l\in\{1,...,p-1\}$ the factor $\lambda(\hat{\zeta}^{l}\cdot V_{l})$ is invertible. Therefore, we will rename $V_{l}=V$ and set $r=\operatorname{rank}(V_{l}).$ By the splitting principle [34, Section 3.2], there exists a smooth projective variety $Z^{\prime}$ and a flat morphism $q:Z^{\prime}\rightarrow Z_{b}$ such that $q^{*}:A^{*}(Z_{b})\rightarrow A^{*}(Z^{\prime})$ is injective and $q^{*}(V)\rightarrow Z^{\prime}$ splits into a direct sum of line bundles, say $q^{*}(V)=\oplus_{b=1}^{r}L_{b}$. In the Chow ring we have that
$\operatorname{ch}_{0}(\lambda(V))=\operatorname{ch}_{0}\left(\prod_{b=1}^{r}\lambda(L_{b})\right)=(1-\hat{\zeta}^{l})^{r}\in S.$ (84)
Further, in the Chow ring we have that
$\operatorname{ch}(\lambda(V))=\operatorname{ch}_{0}(\lambda(V))-(\operatorname{ch}_{0}(\lambda(V))-\operatorname{ch}(\lambda(V))).$ (85)
Since $\operatorname{ch}_{0}(\lambda(V))\in S$ is invertible in the equivariant Chow ring, and since the element $(\operatorname{ch}_{0}(\lambda(V))-\operatorname{ch}(\lambda(V)))$ nilpotent and $A^{*}(Z_{b})$ is a complete graded ring with highest grade equal to $\dim(Z_{b})$, it is clear from (85) that $\operatorname{ch}(\lambda(V))$ is invertible with inverse given by
$\begin{split}&\operatorname{ch}(\lambda(V))^{-1}=\operatorname{ch}_{0}(\lambda(V))^{-1}\left(1-\frac{\operatorname{ch}_{0}(\lambda(V))-\operatorname{ch}(\lambda(V))}{\operatorname{ch}_{0}(\lambda(V))}\right)^{-1}\\ &=\operatorname{ch}_{0}(\lambda(V))^{-1}\sum_{a=0}^{\dim(Z_{b})}\left(\frac{\operatorname{ch}_{0}(\lambda(V))-\operatorname{ch}(\lambda(V))}{\operatorname{ch}_{0}(\lambda(V))}\right)^{a}.\end{split}$ (86)
Since $\operatorname{ch}(\operatorname{rank}(W))=\operatorname{ch}_{0}(W)$ for any equivariant vector bundle $W$ and since (82) is an isomorphism of rings we see that the identities in (81) follows from (84) and (86). Our reason for resorting to the isomorphism (82) is simply that the grading on $A^{*}(Z_{b})$ makes it transparent that $(\operatorname{ch}_{0}(\lambda(V))-\operatorname{ch}(\lambda(V))$ is nilpotent.∎
We now state the Atiyah-Bott Lefshetz formula. Note that by Proposition 2.2 the class $\lambda(N^{*}(Z_{b}))$ is invertible in $S^{-1}\cdot K_{\mu_{p}}^{0}(Z_{b})$ for each $b\in B$.
@@
The proof is a simple computation which combines (LABEL:eq:IdentificationKTheory2) with (LABEL:Lefshetz). The key point is that the identity (LABEL:eq:IdentificationKTheory2) can be generalized to an identity in the $K$-theory ring of $\mathbb{T}\times\mu_{p}$-representations. ∎
We now recall Nielsens’ localization result [62]. As above, we can regard $S$ as a multiplicatively closed subset of $K^{0}_{\mu_{p}}(Z)$ and of $K^{0}_{\mu_{p}}(Z^{\mu_{p}})$.
@@
We now analyze $K_{\mu_{p}}^{0}(Z^{\mu_{p}})$ in more detail. Recall that we have canonical isomorphisms
@@
Recall that for $j\in\{0,...,p-1\}$ we write $\pi_{j}$ for the projection onto the $j$’th summand in (89). We introduce the special notation $W_{0}=\pi_{0}(W).$ We denote the universal localization homomorphism by $L:K^{0}_{\mu_{p}}(Z^{\mu_{p}})\rightarrow S^{-1}\cdot K^{0}_{\mu_{p}}(Z^{\mu_{p}}).$ Recall that the kernel of $L$ is generated by all elements $r$ such that $s\cdot r=0$ for some $s\in S$. Let $Y$ be a formal variable. Recall that $\mathbb{Z}[\mu_{p}]\cong\mathbb{Z}[Y]/(Y^{P}-1)$ and that $Y^{P}-1=(Y-1)\Phi_{p}(Y)$ is the factorization of $Y^{p}-1$ into irreducible factors. As $(1-\hat{\zeta})\in S$, it is therefore clear that we have $).\ker(L)=K^{0}(Z^{\mu_{p}})\otimes_{\mathbb{Z}}(\Phi_{p}(\hat{\zeta}$)). Moreover, the following sequence of abelian groups is split exact
$0\rightarrow\ker(L)\rightarrow K_{\mu_{p}}^{0}(Z^{\mu_{p}})\overset{L}{\rightarrow}\operatorname{Im}(L)\rightarrow 0,$ (90)
@@
[…]
2.1 Application of Localization
Recall now the setup and the notation introduced in Section LABEL:sec:introlocalization. For each $l\in\{0,...,p-1\}$, let $\pi_{l}$ be the projection onto the $l$’th summand of $R_{i,j}$ in (LABEL:decomp).
@@ 2.1 Application of Localization
It will be seen in the proof of Theorem LABEL:ThmMain that the $\tilde{a}_{i,j,l}$ and the $\tilde{b}_{i,j,l}$ are in fact positive integers. Before given the proof of Theorem LABEL:ThmMain we state the following Lemma.
@@
Write $a=\sum_{i=0}^{\infty}a_{i}t^{-i}$ and $b=\sum_{i=0}^{\infty}b_{i}t^{-i}$ with $a_{i},b_{i}\in R$ for all non-negative integers $i$. Without loss of generality, we may assume $a_{0}\not=0$ and $b_{0}\not=0$. We have
@@
In particular, we have that $a_{0}^{p}=b_{0}^{p}$. For each positive integer $l$, let $R^{l}$ denote the degree $2l$ part of $R$ with respect to the cohomological grading. Write $a_{0}=\sum_{l=0}^{d}a_{0,l}$ and $b_{0}=\sum_{l=0}^{d}b_{0,l}$ with $a_{0,l},b_{0,l}\in R^{l}$ for l=0,,dl=0,...,d. Then we have $a_{0,0}=b_{0,0}$ by assumption, and we see that
@@
For each $l=1,..,d$ we get by taking the degree $l$ part of (LABEL:algebra) that
@@
Now notice that any element of $R$ whose projection onto $R^{0}$ is non-zero is not a zero divisor in $R$. In particular $a_{0,0}^{p}=b_{0,0}^{p}$ is not a zero divisor, and therefore we get by induction from (LABEL:algebra2) that $a_{0,l}=b_{0,l}$ for l=0,,dl=0,...,d. Thus $a_{0}=b_{0}$. From (LABEL:algebra0) we deduce that for all non-negative integers $i$ we have that
@@
By induction, and the fact that $a_{0}^{p}$ and $b_{0}^{p}$ are not zero divisors, we obtain from (LABEL:algebra3) that $a_{i}=b_{i}$ for all non-negative integers $i$. Thus $a=b$. ∎
Proof of Theorem LABEL:ThmMain.
We must prove equation (LABEL:eq:Main). The bundle $\mathcal{L}\rightarrow\mathcal{M}$ satisfies the conditions of Proposition 2.1 with the $\mathbb{T}\times\langle f\rangle$-equivariant structure constructed in Section 1. This follows from the fact that $\mathcal{M}$ is a semi-projective variety [44]. Thus we can apply Proposition 2.1. Let $i\in\mathcal{I}$ and $j\in\mathcal{J}_{i}$. To ease notation, we will write $\mathcal{M}_{i,j}^{\mathbb{T}}=\mathcal{M}_{i,j}$ and we will write $T_{i,j}\mathcal{M}$ (resp. $T_{i,j}\mathcal{M}_{i}$) for the restriction of $T\mathcal{M}$ (resp. $T\mathcal{M}_{i}$) to $\mathcal{M}_{i,j}$. By applying Proposition 2.1 we obtain
$\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L}^{k},f)=\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{J}_{i}}(-1)^{\nu_{i,j}}\int_{\mathcal{M}_{i,j}}\Omega(T_{i,j}\mathcal{M})\cdot\operatorname{ch}(\mathcal{L}^{k}_{\mathcal{M}_{i,j}})\cdot\operatorname{Td}(\mathcal{M}_{i,j}).$ (99)
Therefore, proving equation (LABEL:eq:Main) will be sufficient to prove that for all $i\in\mathcal{I}$ and $j\in\mathcal{J}_{i}$ we have that
$(-1)^{\nu_{i,j}}\Omega(T_{i,j}\mathcal{M})$ $=(-1)^{\tilde{\nu}_{i,j}}\varXi_{i}^{*}(\tilde{\chi}_{i,j}),$ (100)
$\operatorname{ch}(\mathcal{L}_{\mid\mathcal{M}_{i,j}})$ $=\hat{\zeta}^{m_{i}}\varXi_{i}^{*}(\exp(\tilde{\omega}_{i,j})).$ (101)
This naturally divides the rest of the proof into two parts. In first, we will prove (100), in the second, we will prove (101).
First part.
We will begin by proving the equation obtained by raising both sides of equation (100) to the $p^{\prime}th$ power. Write $\mathbb{U}=\operatorname{End}_{0}(\operatorname{\mathbbm{E}})$, and define the equivariant class
@@
This can be seen for instance by considering a standard long exact sequence for hypercohomology groups of two-term chain complexes (see [42, §2.2.3]). Consider the natural projection onto the first factor $\pi^{\prime}:\mathcal{M}_{i,j}\times X^{f}\rightarrow\mathcal{M}_{i,j}$. Note that we have $(\mathcal{M}_{i,j}\times X)^{f}=\mathcal{M}_{i,j}\times X^{f}$. Consider the inclusion $\iota:\mathcal{M}_{i,j}\times X^{f}\rightarrow\mathcal{M}_{i,j}\times X$. Clearly $\pi\circ\iota=\pi^{\prime}.$ By applying Nielsen’s localization theorem to the right hand side of quation (103), we obtain the following identity in $S^{-1}\cdot K^{0}_{\mathbb{T}\times\mu_{p}}(\mathcal{M}_{i,j})$
@@
However, the $\mathbb{T}\times\mu_{p}$-equivariant $K$-theory operation $s$ of taking the sum of all symmetric powers is not well-defined in $S^{-1}\cdot K^{0}_{\mathbb{T}\times\mu_{p}}(\mathcal{M}_{i,j})$, and we need to extract an identity in the ring $K_{\mathbb{T}\times\mu_{p}}^{0}(\mathcal{M}_{i,j})$. For all u=1,,p1u=1,...,p-1, we have the following identity in $S^{-1}\cdot\mathbb{Z}[\mu_{p}]$